OpenStudy (anonymous):

Determine the solution set of (x - 3)2 = 49. {6, -8} {4, 10} {10, -4}

OpenStudy (anonymous):

(x - 3)^2 = 49 (the two is an exponent)

OpenStudy (solomonzelman):

\(\large\color{black}{(x - 3)^2 = 49 }\) re-write the right side, \(\large\color{black}{(x - 3)^2 = 7^2 }\) square both sides, \(\large\color{black}{x - 3 = \pm 7 }\) from here, you should be able to finish...

OpenStudy (solomonzelman):

please next time use the math section for math questions. (Click more subject, and search for "mathematics")

OpenStudy (anonymous):

The answer is (10,-4)